OpenStudy (pgxgirl103):

Solve for x? √ x+3 =2

OpenStudy (anonymous):

square both sides

OpenStudy (anonymous):

Assuming that the prob. is √(x+3) = 2 (NOT, √x + 3 = 2 )

OpenStudy (pgxgirl103):

yes sorry and square x and 3?

OpenStudy (anonymous):

A quick example of squaring both sides... √(b-4)=2 [ √(b-4) ] ² = 2 ² b - 4 = 4 b = 8 OKAY? Can you do your problem similarly ?

OpenStudy (anonymous):

Squaring both sides, means to raise each (entire) side (of the equation) to the second power.

OpenStudy (pgxgirl103):

√ (x+3)^2=2^2 is that right?

OpenStudy (anonymous):

Yes !!!

OpenStudy (anonymous):

What are you getting :) ?

OpenStudy (pgxgirl103):

I don't know what to do next.

OpenStudy (anonymous):

What will the equation look like after you square both sides (you started it well, go on !)

OpenStudy (pgxgirl103):

3 squared is 9 and 2 squared is 4.

OpenStudy (anonymous):

√ (x+3) ² =2 ² √ (x+3) ² = x+3 2 ² = 4 So, ... x+3=4 and the last step ?

OpenStudy (pgxgirl103):

you do 4-3? which equals 1 so x=1?

OpenStudy (pgxgirl103):

@student_basil

OpenStudy (anonymous):

yes, you got it, correct, x=1 .

OpenStudy (pgxgirl103):

thanks!